OpenStudy (anonymous):

Solve on the interval [0,2pi)

OpenStudy (anonymous):

sin 2x = 2 sin x cos x solve sin 2x = - sin x sin 2 x + sin x = 0 2 sin x cos x + sin x = 0 sin x( 2 cos x + 1) = 0 sin x = 0 x = n pi : n = 0, 1, 2 2 cos x + 1 = 0 cos x = -1/2 x = 2 pi/3 and 4 pi/ 3 ans: x = 0, 2 pi/3, pi, 4 pi/3, 2 pi